How many C atoms are present in the sample of C3H8 with mass 4.41 g? Avogadro’s number is 6.022 × 1023. Answer using scientific notation and to three significant digits.
Answer #1

Molar mass of C3H8,

MM = 3*MM(C) + 8*MM(H)

= 3*12.01 + 8*1.008

= 44.094 g/mol

mass(C3H8)= 4.41 g

use:

number of mol of C3H8,

n = mass of C3H8/molar mass of C3H8

=(4.41 g)/(44.094 g/mol)

= 0.100 mol

use:

number of molecules = number of mol * Avogadro’s number

number of molecules = 0.1 * 6.022*10^23 molecules

number of molecules = 6.023*10^22 molecules

Total number of C in 1 molecule = 3

use:

number of atoms of C= number of molecules * Total number of C atoms in 1 molecule

= 6.023*10^22 * 3

= 1.81*10^23

Answer: 1.81*10^23